OpenStudy (anonymous):

A 1.4 kg ball is dropped from an unknown height, h, and hits the ground with a velocity of 15.0 m/s and rebounds with a velocity of 11.0 m/s. If the ball is in contact with the ground for 1/35 of a second find the average force the ground applies to the ball.

OpenStudy (anonymous):

[F =m( (v-u) \t)] where F=force v=final velocity u=initial velocity m=mass t=time. Using this formula impulse we get a answer of -196N

OpenStudy (anonymous):

I tried that but it wasn't right D: the answer is suppose to be 1288N, thanks for trying though:)

OpenStudy (anonymous):

yes the final velocity should be considered as "-11" now the answer will be 1274 the reason for taking final velocity as negative is that it is in the opposite direction as the initial velocity

OpenStudy (anonymous):

average force = change of momentum of the ball impact = Ns = mv - mu N = (mv-mu)/ s = (1.4*15 --1.4*11) / 1/35 = 1274 newtons

OpenStudy (anonymous):

THANKS SO MUCH GUYS! This really helped!!!